OpenStudy (anonymous):

How do you simplify 1/4y - 2 1/4y +1/2(4 - 2y)

OpenStudy (anonymous):

2-y+5/y

OpenStudy (anonymous):

how did you get that

OpenStudy (anonymous):

for the first and second terms \[1\div4y-21\div4y=-20\div4y\]

OpenStudy (anonymous):

for the third term \[1\div2\left( 4-2y \right)=2-y\]

OpenStudy (anonymous):

then the sum \[-20\div4y +2-y\]

OpenStudy (anonymous):

thanks